Minutes of the Booster Commissioning Working Group held on 3rd of July 2008

Present: K. Hanke, W. Weterings, B. Mikulec.


  1. Communications
  2. Follow-up of open actions
  3. Planning of PSB commissioning with Linac4
  4. AOB

2. Follow-up of open actions:

  • Kicker magnets in LBE line: With the current magnet parameters it could be clarified that the magnet is really a kicker. Depending on the simulation results for the LBE line and the new specifications, the responsibility for the new magnet will be determined.

  • Early removal of lead distributor: S. de Man has been contacted to give his opinion whether the removal of the lead distributor could be advanced to the 2011/2012 shutdown. Awaiting his answer.

3. Planning of PSB commissioning with Linac4:

The planning for the PSB commissioning has been modified corresponding to the discussion last meeting and accepted by M. Vretenar. 1 month cool-down time is marked as a separate task, 3 months (plus the Christmas holiday period) have been attributed to the modification of the PSB injection period and 3 months to PSB beam commissioning with Linac4.
Cabling and power converter modifications can already be started in parallel to the cool-down period.

L4_test_installation_commissioning_v4.mpp: updated installation and commissioning planning

To this time the period for PS and SPS commissioning has to be added to prepare the LHC beams originating from the new operation with Linac4; tentatively 2 weeks have been added per machine.

Going through the different tasks it is unclear from which point onwards no operation with Linac2 will be possible anymore. Where is the connection made between the new transfer line and the old Linac2 transfer line? A hole has to be made to connect the 2 tunnels - is this hidden e.g. in the task 'Upgrade Linac2 line' (shutdown 2011/2012)? Moreover LT.BHZ20 will have to be changed for Linac4 operation; how long does this take? Can this time be included in the task 'Commissioning PIMS'?
All together the shutdown will amount to about 10 months.

4. AOB:

It has been decided to wait for detailed error studies by ABP before fixing the parameters in the BI line (distributor, septum, BI.DVT), but after a recent presentation of M. Eshraqi (see Different optics for the transfer line) it looks promising that the old scheme with the lower kicker strength can be kept.

-- BettinaMikulec - 03 Jul 2008

Topic attachments
I Attachment History Action Size Date Who Comment
Unknown file formatmpp L4_test_installation_commissioning_v4.mpp r1 manage 140.0 K 2008-08-22 - 09:01 KlausHanke updated installation and commissioning planning
Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2008-08-22 - KlausHanke
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    SPL All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright